Find three consecutive odd integers such that twice the second is 5 more than the last ?
Answer by venugopalramana(3286)

Find three consecutive odd integers such that twice the second is 5 more than the last ?Find three consecutive odd integers such that twice the second
is 5 more than the last ?
ANY INTEGER IS N
ANY EVEN INTEGER IS 2N
ANY ODD INTEGER IS 2N-1
3 CONSECUTIVE ODD INTEGERS ARE
2N-1,2N-1+2=2N+1,2N-1+2+2=2N+3
TWICE THE SECOND = 2(2N+1)=4N+2
5 MORE THAN THE LAST = 2N+3+5=2N+8....HENCE
4N+2=2N+8
4N-2N=8-2=6
2N=6
N=3.
HENCE THE 3 NUMBERS ARE
2N-1=2*3-1=5
2N+1=2*3+1=7
2N+3=2*3+3=9